# Hot, Photoionized X-ray Gas in Two Luminous Type 2 Quasars: Chandra-HST Evidence for a Wind-Driven Sequence

### v1

@@ 7 An Evolutionary Path for Quasars
The QSO2s analyzed by fischer_hubble_2018 span a wide range of kinematic and morphological properties. Their sample includes twelve of the fifteen most luminous type 2 quasars in the reyes_space_2008 catalog with . All targets have erg s-1, placing them in the brightest of the QSO2 population from reyes_space_2008. The corresponding black hole masses lie in the range – (e.g., fischer_hubble_2018). Most sources are radio-quiet, with a minority classified as intermediately radio-loud, indicating that their observed outflows are unlikely to be jet-driven.
@@
All twelve objects host [O III]-emitting NLRs that extend beyond 1 kpc, yet only two show outflows that propagate on similarly large scales. fischer_hubble_2018 argued that the distribution of circumnuclear gas near the nucleus may regulate how far ionizing radiation and winds can penetrate into the NLR: in morphologically compact systems, dense inner material can absorb a significant fraction of the AGN ionizing flux, confining outflows to smaller radii. Expanding on this picture, trindade_falcao_hubble_2021 proposed that X-ray winds may progressively entrain this circumnuclear gas, eventually enabling a transition into a large-scale blow-out phase.
@@
The differing propagation timescales of ionizing radiation and mechanical feedback naturally account for systems with extended ionized [O III] emission but no large-scale winds. A wind traveling at km s-1 requires yr to cross 1 kpc, whereas the light-crossing time to the same distance is only yr. Consequently, photoionization can illuminate kiloparsec-scale [O III] structures long before mechanical winds have propagated to comparable radii. This evolutionary framework, summarized schematically in Figure 8, links the diversity of morphologies and kinematics observed by fischer_hubble_2018 to different stages along a continuum sequence of circumnuclear clearing and feedback coupling.
Our Chandra observations support this proposed scenario. The X-ray and [O III] properties of 2MASX J130038 and FIRST J120041 suggest that they occupy different locations along the early portion of this evolutionary sequence. In 2MASX J130038, the X-ray emission is extended but only weakly correlated with the [O III] morphology, while the optical kinematics remain dominated by ordered rotation with localized turbulence near the nucleus. Combined with the tentative detection of a blueshifted Fe XXVI line at km s-1 (Section 6.2), these features point to a system in the onset of a blow-out stage (Stage II; see Figure 8), in which a nascent hot wind is beginning to leak out of the nucleus and interact with the inner circumnuclear medium.
@@
Taken together, these quasars may trace the transition from a kinematically compact, blow-out configuration (Stage II) to the onset of a more extended, early post–blow-out state (Stage III). For clarity, in Figure 8 we illustrate this sequence through four representative “stages,” while Table 6 summarizes how the twelve QSO2s studied by fischer_hubble_2018 may populate these stages. These are labeled as follows:
- Stage I: Pre-Blow-Out
@@
This initial phase represents heavily enshrouded systems in which both the X-ray and [O III] emission are morphologically compact and centrally concentrated. Ionizing radiation and winds remain trapped by dense circumnuclear gas, preventing large-scale outflows to eventually form. In the fischer_hubble_2018 sample, examples may include 2MASX J11001238, 2MASX J14054117, and 2MASX J17135038, which all show compact [O III] morphologies ( kpc; fischer_hubble_2018) but broad nuclear line widths (FWHM1000 km s-1; fischer_hubble_2018) indicative of turbulent but spatially confined gas in the inner region.
- Stage II: Blow-Out
@@
Systems where the AGN ionizing radiation begins to escape through low-density channels, ionizing more extended structures but still lacking large-scale outflows. 2MASX J130038 exemplifies this behavior: its [O III] emission extends to kpc, yet the gas is largely rotational, with only localized disturbances within 160 pc of the nucleus (fischer_hubble_2018). The X-rays are morphologically centrally peaked and weakly correlated with [O III] emission (Figure 4), implying that the hot and cooler [O III] phases remain decoupled. The tentative identification of Fe XXVI emission at 7.14 keV ( km s-1) could point to the onset of a nascent hot wind beginning to interact with the inner circumnuclear gas. Similar Stage II behavior is seen in other systems such as 2MASX J07594101, 2MASX J080252, SDSS J115245, B2 1435, Mrk 477, and 2MASX J16531506, where [O III] ionized emission is extended (3 kpc; fischer_hubble_2018) but the kinematics remain dominated by broad line widths and low velocity centroids (FWHM 600 km s-1, 600 km s-1; fischer_hubble_2018). The large X-ray gas reservoir found in 2MASX J130038 likely reflects the accumulation of material during the first evolutionary phases, when dense circumnuclear gas remained confined close to the AGN. As this gas lies nearer to the ionizing source, it reaches higher ionization states, naturally producing a more massive X-ray–emitting component at the onset of blow-out.
- Stage III: Early Post-Blow-Out
@@
As the wind begins to break through the circumnuclear material, the X-ray outflow expands, producing outflowing [O III] kinematics and the spatially correlated X-ray and optical emission we observe. FIRST J120041 shows bright [O III] emission extending to kpc, with multi-component, high-velocity profiles (-750 km s-1) and broad line widths (FWHM 1700 km s-1; fischer_hubble_2018). The soft X-ray emission is clumpy and structurally complex, showing a strong morphological correspondence with the [O III] gas (Figure 3). Peaks in the soft X-ray brightness are co-spatial with regions of outflowing [O III] gas (Figure 6), suggesting that the X-ray wind is now entraining and driving the optical phase. These properties are consistent with a system in which a blow-out has recently occurred, marking the transition from a confined to an expanding feedback regime.
- Stage IV: Late Post-Blow-Out
@@
The nuclear X-ray winds have now cleared most of the circumnuclear material, driving powerful, large-scale outflows. Mrk 34 exemplifies this phase: its [O III] emission extends in a bicone to kpc, showing broad, high-velocity components ( km s-1, FWHM 1700 km s-1; fischer_hubble_2018). It hosts the highest [O III] mass outflow rate and kinetic luminosity in the sample, suggesting that the dense circumnuclear medium has been largely expelled, allowing the AGN to efficiently couple its energy to the host ISM and sustain galaxy-scale feedback (trindade_falcao_hubble_2021-1).
@@
Taken together, these stages outline a coherent evolutionary pathway for quasars. The sequence from pre-blow-out to late post-blow-out phases reflects the gradual coupling of the X-ray and [O III] ionized gas phases as the AGN clears its environment, transitioning from compact, enshrouded systems to those exhibiting large-scale, energetically dominant outflows. A subsequent stage in this sequence may be represented by (unobscured) type 1 quasars (QSO1s), in which the circumnuclear gas has been completely cleared, giving the AGN a “naked” appearance, an unobscured view of the central engine once feedback has effectively evacuated the surrounding medium.
@@
Supporting this scenario, trindade_falcao_hubble_2024 found that the sizes of [O III]-emitting NLRs in type 1 AGNs (spanning erg s-1) scale with luminosity as , a significantly steeper relation than that found for type 2 AGNs over the same luminosity range, . The combined type 1 + type 2 population follows an intermediate slope of . For the subset of seven QSO1s, trindade_falcao_hubble_2024 additionally found significantly extended radial mass profiles, with the bulk of the [O III]–emitting gas located at radii kpc. Relative to the twelve QSO2s in fischer_hubble_2018, these QSO1s have NLRs larger by roughly a factor of three yet contain smaller total [O III] gas masses, consistent with systems in which feedback has already cleared much of the circumnuclear medium and redistributed the remaining ionized gas to larger scales.
@@
| 2MASX J07594101 | Extended, conical | 3.2 | 0.67 | 315 | 0.67 | 1680 | II – Blow-Out |
| 2MASX J08025293 | Extended, biconical | 3.6 | 0.57 | 700 | 0.89 | 870 | II – Blow-Out |
| Mrk 34 | Extended, biconical | 2.2 | 1.89 | 1100 | 1.89 | 580 | IV – Late Post-Blow-Out |
| 2MASX J11001238 | Compact | 3.0 | 0.68 | 350 | 1.51 | 1780 | I – Pre-Blow-Out |
| SDSS J115245 | Extended, biconical | 3.6 | 0.15 | 300 | 1.23 | 360 | II – Blow-Out |
| FIRST J120041 | Extended, biconical | 6.1 | 1.07 | 450 | 1.59 | 720 | III – Early Post-Blow-Out |
| 2MASX J130038 | Extended, biconical | 4.7 | 0.16 | 100 | 0.16 | 580 | II – Blow-Out |
| 2MASX J14054117 | Compact | 1.0 | 0.33 | 100 | 0.94 | 760 | I – Pre-Blow-Out |
| B2 1435 | Extended, biconical | 3.8 | 0.20 | 200 | 1.74 | 630 | II – Blow-Out |
| Mrk 477 | Extended, conical | 2.7 | 0.54 | 500 | 0.90 | 2040 | II – Blow-Out |
| 2MASX J16531506 | Extended, biconical | 6.50 | 0.57 | 110 | 1.23 | 510 | II – Blow-Out |
| 2MASX J17135038 | Compact | 1.8 | 0.65 | 160 | 0.92 | 1660 | I – Pre-Blow-Out |
8 Summary and Conclusions
We present new Chandra/ACIS-S imaging spectroscopy of two luminous QSO2s from the fischer_hubble_2018 sample, FIRST J120041.4+314745 and 2MASX J13003807+5454367, and compare their X-ray properties with HST [O III] morphologies and kinematics. Our analysis combines PSF modeling and subtraction (Section 4.1), broad- and narrow-band imaging (Section 4.2), phenomenological spectral modeling (Section 3.1), and physically motivated photoionization modeling using Cloudy (Section 3.2, and Section 5). The main results are as follows:
@@ 8 Summary and Conclusions
Both quasars show extended X-ray emission, but with different levels of correlation to the optical phase. In FIRST J120041, the X-rays form clumpy, kiloparsec-scale structures that closely trace the [O III] bicone (Figure 3); peaks in soft X-ray surface brightness coincide with the regions of highest [O III] velocities (600-750 km s-1) and broadest line widths (FWHM up to 1700 km s-1; Figure 6). In contrast, 2MASX J130038 shows centrally concentrated X-ray emission with weak morphological correspondence to the [O III] morphology (Figure 4) and no clear spatial correlation with the largely rotational [O III] kinematics (Figure 7).
@@
The inferred hot-gas reservoirs are massive. Within apertures (2 kpc for FIRST J120041 and 1.6 kpc for 2MASX J130038), we estimate and , respectively, based on photoionization modeling (Table 5). Relative to the outflowing [O III] gas (trindade_falcao_hubble_2021-1), the X-ray phase is roughly 4 times more massive (volume-normalized) in FIRST J120041 and 16 times more massive in 2MASX J130038. These results indicate that the hot, photoionized phase can dominate both the mass and energy budgets of the ionized ISM on kiloparsec scales.
@@
2MASX J130038 shows evidence for a nuclear hot wind. An emission feature at keV is consistent with blueshifted Fe XXVI Ly ( km s-1), suggesting a highly-ionized component near the nucleus. This emission may trace a nascent hot wind beginning to emerge within the inner few hundred parsecs (Section 6.2).
@@
A continuous evolutionary sequence may connect the two quasars and the broader QSO2 population. Integrating the X-ray and optical morphologies, kinematics, and gas masses, we propose that luminous type 2 quasars evolve along a continuum sequence regulated by the gradual clearing of circumnuclear material by AGN radiation and winds (Section 7; Figure 8). In this framework, 2MASX J130038 corresponds to an earlier phase where a weakly coupled, emerging X-ray wind begins to interact with the ISM (analogous to a “blow-out” configuration), while FIRST J120041 represents a relatively more developed stage where the hot wind actively entrains and drives the [O III] gas (an “early post-blow-out” configuration). More compact, enshrouded systems such as 2MASX J11001238 or 2MASX J17135038 may represent the most buried, confined phase of this sequence, “pre-blow-out”, whereas sources like Mrk 34 illustrate the opposite extreme, a “late post-blow-out” characterized by powerful, large-scale outflows and efficient AGN-ISM coupling (Table 6).
@@
Finally, the broader distribution of the fischer_hubble_2018 QSO2 sample across this sequence of clearing and coupling implies that luminous obscured quasars may represent different points along a single evolutionary sequence driven by AGN feedback, likely culminating in unobscured type 1 quasars (trindade_falcao_hubble_2024), once the circumnuclear medium has been fully dispersed and the central engine is directly exposed.

### v2

@@ VII An Evolutionary Path for Quasars
The QSO2s analyzed by Fischer et al. (2018) span a wide range of kinematic and morphological properties. Their sample includes twelve of the fifteen most luminous type 2 quasars in the Reyes et al. (2008) catalog with . All targets have erg s-1, placing them in the brightest of the QSO2 population from Reyes et al. (2008). The corresponding black hole masses lie in the range – (e.g., Fischer et al., 2018). Most sources are radio-quiet, with a minority classified as intermediately radio-loud, indicating that their observed outflows are unlikely to be jet-driven. Both quasars analyzed in this work, FIRST J120041 and 2MASX J130038, are classified as radio-quiet objects, and therefore belong to the radio-quiet majority of the sample.
@@
All twelve objects host [O III]-emitting NLRs that extend beyond 1 kpc, yet only two show outflows that propagate on similarly large scales. Fischer et al. (2018) argued that the distribution of circumnuclear gas near the nucleus may regulate how far ionizing radiation and winds can penetrate into the NLR: in morphologically compact systems, dense inner material can absorb a significant fraction of the AGN ionizing flux, confining outflows to smaller radii. Building on this picture, Trindade Falcão et al. (2021a) suggested that inner X-ray winds may progressively entrain circumnuclear gas, an idea that has also been explored in theoretical models of AGN feedback (e.g., Costa et al., 2014) and supported by observational studies tracing multi-phase outflows through X-ray absorption features (e.g., Serafinelli et al., 2019; Xu et al., 2024). In this framework, continued entrainment may eventually enable a transition into a large-scale blowout phase (Trindade Falcão et al., 2021a).
@@
The differing propagation timescales of ionizing radiation and mechanical feedback naturally account for systems with extended ionized [O III] emission but no large-scale winds. A wind traveling at km s-1 requires yr to cross 1 kpc, whereas the light-crossing time to the same distance is only yr. Consequently, photoionization can illuminate kiloparsec-scale [O III] structures long before mechanical winds have propagated to comparable radii. This evolutionary framework, summarized schematically in Figure 8, links the diversity of morphologies and kinematics observed by Fischer et al. (2018) to different stages along a continuum sequence of circumnuclear clearing and feedback coupling.
Although this framework is developed using a low-redshift QSO2 sample, it is not intended to be tied to fixed luminosities or physical scales. Instead, the proposed stages describe relative evolutionary stages regulated by obscuration and feedback coupling. At higher redshifts and luminosities, such as in Hot Dust-Obscured Galaxies (Hot DOGs; see, e.g., Eisenhardt et al. 2012; Assef et al. 2015), the same sequence may operate at larger energies and on shorter timescales. In this context, Hot DOGs can be interpreted as extreme analogs of the late pre-blowout and blowout phases, in which systems remain heavily obscured while already exhibiting powerful ionized outflows traced by [O III], even as direct X-ray signatures of the inner wind are attenuated by large column densities (e.g., Costa et al., 2014; Serafinelli et al., 2019; Xu et al., 2024; Villani et al., 2026). Objects that fully reach a late post-blowout phase would likely no longer satisfy Hot DOG selection criteria.
Our Chandra observations support this proposed scenario (see Figure 8). The X-ray and [O III] properties of 2MASX J130038 and FIRST J120041 suggest that they occupy different locations along the early portion of this evolutionary sequence. In 2MASX J130038, the X-ray emission is extended but only weakly correlated with the [O III] morphology, while the optical kinematics remain dominated by ordered rotation with localized turbulence near the nucleus. Combined with the tentative detection of a blueshifted Fe XXVI line at km s-1 (Section VI.2), these features point to a system in the onset of a blowout stage (Stage II; see Figure 8), in which a nascent hot wind is beginning to leak out of the nucleus and interact with the inner circumnuclear medium.
@@
Taken together, these quasars may trace the transition from a kinematically compact, blowout configuration (Stage II) to the onset of a more extended, early post–blowout state (Stage III). We emphasize that the characteristic spatial scales quoted below (e.g., kpc for Stage I or -6 kpc for later stages) refer specifically to the objects in the Fischer et al. (2018) sample discussed here. These values are intended as empirical descriptors rather than universal thresholds, and are expected to depend on intrinsic AGN properties such as black hole mass and luminosity. For clarity, in Figure 8 we illustrate this sequence through four representative “stages,” while Table 6 summarizes how the twelve QSO2s studied by Fischer et al. (2018) may populate these stages. These are labeled as follows:
- Stage I: Pre-Blowout
@@
This initial phase represents heavily enshrouded systems in which both the X-ray and [O III] emission are morphologically compact and centrally concentrated. Ionizing radiation and winds remain trapped by dense circumnuclear gas, preventing large-scale outflows to eventually form. In the Fischer et al. (2018) sample, examples may include 2MASX J11001238, 2MASX J14054117, and 2MASX J17135038, which all show compact [O III] morphologies ( kpc; Fischer et al. 2018) but broad nuclear line widths (FWHM1000 km s-1; Fischer et al. 2018) indicative of turbulent but spatially confined gas in the inner region.
- Stage II: Blowout
@@
Systems where the AGN ionizing radiation begins to escape through low-density channels, ionizing more extended structures but still lacking large-scale outflows. 2MASX J130038 exemplifies this behavior: its [O III] emission extends to kpc, yet the gas is largely rotational, with only localized disturbances within 160 pc of the nucleus (Fischer et al., 2018). The X-rays are morphologically centrally peaked and weakly correlated with [O III] emission (Figure 4), implying that the hot X-ray and cooler [O III] phases remain decoupled. The tentative identification of Fe XXVI emission at 7.14 keV ( km s-1; Section VI.2) could point to the onset of a nascent hot wind beginning to interact with the inner circumnuclear gas, but is not required for the Stage II classification. The evolutionary interpretation is instead driven primarily by the spatial decoupling between the X-ray and [O III] emission and by the lack of large-scale outflow kinematics. Similar Stage II behavior is seen in other systems such as 2MASX J07594101, 2MASX J080252, SDSS J115245, B2 1435, Mrk 477, and 2MASX J16531506, where [O III] ionized emission is extended (3 kpc; Fischer et al. 2018) but the kinematics remain dominated by broad line widths and low velocity centroids (FWHM 600 km s-1, 600 km s-1; Fischer et al. 2018). The large X-ray gas reservoir found in 2MASX J130038 likely reflects the accumulation of material during the first evolutionary phases, when dense circumnuclear gas remained confined close to the AGN. As this gas lies nearer to the ionizing source, it reaches higher ionization states, naturally producing a more massive X-ray-emitting component at the onset of blowout.
- Stage III: Early Post-Blowout
@@
As the wind begins to break through the circumnuclear material, the X-ray outflow expands, producing outflowing [O III] kinematics and the spatially correlated X-ray and optical emission we observe. FIRST J120041 shows bright [O III] emission extending to kpc, with multi-component, high-velocity profiles (-750 km s-1) and broad line widths (FWHM 1700 km s-1; Fischer et al. 2018). The soft X-ray emission is clumpy and structurally complex, showing a strong morphological correspondence with the [O III] gas (Figure 3). Peaks in the soft X-ray brightness are co-spatial with regions of outflowing [O III] gas (Figure 6), suggesting that the X-ray wind is now entraining and driving the optical phase. These properties are consistent with a system in which a blowout has recently occurred, marking the transition from a confined to an expanding feedback regime.
- Stage IV: Late Post-Blowout
@@
The nuclear X-ray winds have now cleared most of the circumnuclear material, driving powerful, large-scale outflows. Mrk 34 exemplifies this phase: its [O III] emission extends in a bicone to kpc, showing broad, high-velocity components ( km s-1, FWHM 1700 km s-1; Fischer et al. 2018). It hosts the highest [O III] mass outflow rate and kinetic luminosity in the sample, suggesting that the dense circumnuclear medium has been largely expelled, allowing the AGN to efficiently couple its energy to the host ISM and sustain galaxy-scale feedback (Trindade Falcão et al., 2021b).
@@
Taken together, these stages outline a coherent evolutionary pathway for quasars. The sequence from pre-blowout to late post-blowout phases reflects the gradual coupling of the X-ray and [O III] ionized gas phases as the AGN clears its environment, transitioning from compact, enshrouded systems to those exhibiting large-scale, energetically dominant outflows. A subsequent stage in this sequence may be represented by (unobscured) type 1 quasars (QSO1s), in which the circumnuclear gas has been completely cleared, giving the AGN a “naked” appearance, an unobscured view of the central engine once feedback has effectively evacuated the surrounding medium.
@@
Supporting this scenario, Trindade Falcão et al. (2024) found that the sizes of [O III]-emitting NLRs in type 1 AGNs (spanning erg s-1) scale with luminosity as , a significantly steeper relation than that found for type 2 AGNs over the same luminosity range, . The combined type 1 + type 2 population follows an intermediate slope of . For the subset of seven QSO1s, Trindade Falcão et al. (2024) additionally found significantly extended radial mass profiles, with the bulk of the [O III]-emitting gas located at radii kpc. Relative to the twelve QSO2s in Fischer et al. (2018), these QSO1s have NLRs larger by roughly a factor of three, yet contain smaller total [O III] gas masses, consistent with systems in which feedback has already cleared much of the circumnuclear medium and redistributed the remaining ionized gas to larger scales.
@@
| 2MASX J07594101 | Extended, conical | 3.2 | 0.67 | 315 | 0.67 | 1680 | II – Blowout |
| 2MASX J08025293 | Extended, biconical | 3.6 | 0.57 | 700 | 0.89 | 870 | II – Blowout |
| Mrk 34 | Extended, biconical | 2.2 | 1.89 | 1100 | 1.89 | 580 | IV – Late Post-Blowout |
| 2MASX J11001238 | Compact | 3.0 | 0.68 | 350 | 1.51 | 1780 | I – Pre-Blowout |
| SDSS J115245 | Extended, biconical | 3.6 | 0.15 | 300 | 1.23 | 360 | II – Blowout |
| FIRST J120041 | Extended, biconical | 6.1 | 1.07 | 450 | 1.59 | 720 | III – Early Post-Blowout |
| 2MASX J130038 | Extended, biconical | 4.7 | 0.16 | 100 | 0.16 | 580 | II – Blowout |
| 2MASX J14054117 | Compact | 1.0 | 0.33 | 100 | 0.94 | 760 | I – Pre-Blowout |
| B2 1435 | Extended, biconical | 3.8 | 0.20 | 200 | 1.74 | 630 | II – Blowout |
| Mrk 477 | Extended, conical | 2.7 | 0.54 | 500 | 0.90 | 2040 | II – Blowout |
| 2MASX J16531506 | Extended, biconical | 6.50 | 0.57 | 110 | 1.23 | 510 | II – Blowout |
| 2MASX J17135038 | Compact | 1.8 | 0.65 | 160 | 0.92 | 1660 | I – Pre-Blowout |
VIII Summary and Conclusions
We present new Chandra/ACIS-S imaging spectroscopy of two luminous QSO2s from the Fischer et al. (2018) sample, FIRST J120041.4+314745 and 2MASX J13003807+5454367, and compare their X-ray properties with HST [O III] morphologies and kinematics. Our analysis combines PSF modeling and subtraction (Section IV.1), broad- and narrow-band imaging (Section IV.2), phenomenological spectral modeling (Section III.1), and physically motivated photoionization modeling using Cloudy (Section III.2, and Section V). The main results are as follows:
@@ VIII Summary and Conclusions
Both quasars show extended X-ray emission, but with different levels of correlation to the optical phase. In FIRST J120041, the X-rays form clumpy, kiloparsec-scale structures that closely trace the [O III] bicone (Figure 3); peaks in soft X-ray surface brightness coincide with the regions of highest [O III] velocities (600-750 km s-1) and broadest line widths (up to 1700 km s-1; Figure 6). In contrast, 2MASX J130038 shows centrally concentrated X-ray emission with weak morphological correspondence to the [O III] morphology (Figure 4) and no clear spatial correlation with the largely rotational [O III] kinematics (Figure 7).
@@
The inferred hot-gas reservoirs are massive. Within apertures (2 kpc for FIRST J120041 and 1.6 kpc for 2MASX J130038), we estimate and , respectively, based on photoionization modeling (Table 5). Relative to the outflowing [O III] gas (Trindade Falcão et al., 2021b), the X-ray phase is roughly 4 times more massive (volume-normalized) in FIRST J120041 and 16 times more massive in 2MASX J130038. These results indicate that the hot, photoionized phase can dominate both the mass and energy budgets of the ionized ISM on kiloparsec scales.
@@
2MASX J130038 shows a marginal evidence high-energy emission feature consistent with a nuclear hot wind. An emission feature at keV improves the spectral fit by Cstat = 10 for two additional free parameters, corresponding to a 2.7 detection. If interpreted as blueshifted Fe XXVI Ly ( km s-1), this feature would be consistent with highly ionized gas in the immediate nuclear region. Given its modest statistical significance, we regard this feature as suggestive rather than conclusive evidence for a nascent hot wind (Section VI.2).
@@
A continuous evolutionary sequence may connect the two quasars and the broader QSO2 population. Integrating the X-ray and optical morphologies, kinematics, and gas masses, we propose that luminous type 2 quasars evolve along a continuum sequence regulated by the gradual clearing of circumnuclear material by AGN radiation and winds (Section VII; Figure 8). In this framework, 2MASX J130038 corresponds to an earlier phase where a weakly coupled, emerging X-ray wind begins to interact with the ISM (analogous to a “blowout” configuration), while FIRST J120041 represents a relatively more developed stage where the hot wind actively entrains and drives the [O III] gas (an “early post-blowout” configuration). More compact, enshrouded systems such as 2MASX J11001238 or 2MASX J17135038 may represent the most buried, confined phase of this sequence, “pre-blowout”, whereas sources like Mrk 34 illustrate the opposite extreme, a “late post-blowout” characterized by powerful, large-scale outflows and efficient AGN-ISM coupling (Table 6).
@@
Finally, the broader distribution of the Fischer et al. (2018) QSO2 sample across this sequence of clearing and coupling implies that luminous obscured quasars may represent different points along a single evolutionary sequence driven by AGN feedback, likely culminating in unobscured type 1 quasars (Trindade Falcão et al., 2024), once the circumnuclear medium has been fully dispersed and the central engine is directly exposed.